# Llama-3.1-Sherkala-8B-Chat: An Open Large Language Model for Kazakh

### v1

@@ 4.1.1 General Instruction-Tuning Data
We curate a set of English instructions in a prompt–response pair format spanning a comprehensive range of NLP and other generative tasks. We use common datasets such as TuluV3 (450,817 samples) [40], and also some of our data is a subset of the instruction-tuning data used for building Jais [72], and hence, is a combination of several publicly available datasets. We also add a significant amount of proprietary data that we gather from relevant internal sources.
Russian
@@
We incorporate 263K instruction-response pairs from the Grandmaster-PRO-MAX [^4] https://huggingface.co/datasets/Vikhrmodels/GrandMaster-PRO-MAX dataset, a high-quality Russian instruction-tuning corpus. This dataset covers a wide range of topics, including mathematics, coding problems, and role-playing scenarios, providing diverse linguistic and task-specific coverage for instruction fine-tuning.
4.1.2 Safety Data
@@
We developed a comprehensive safety prompt collection process specifically tailored for Kazakh model training, covering eight types of attacks and over 100 detailed safety categories. [^5] In the current released version, we randomly sampled 20K data for SFT. To ensure high-quality data, a team of five expert annotators initially crafted “seed prompts” for direct attack alignment based on our previous work [82], resulting in approximately 1,200 annotated examples focused both on general and Kazakhstan-specific scenarios. Building on this foundation, the expert team guided a 20-member outsourced annotation team, leveraging LLMs, to generate an additional 50K attack prompts, ensuring diversity, linguistic relevance, and thorough coverage for Kazakh.
We enrich the set of direct attack prompts in SFT data with a collection of adversarial prompt attack methods. Following [41], we adopt eight adversarial prompt attack methods to construct the SFT data. These methods target the following abilities of LLMs: in-context learning, auto-regressiveness, instruction following, and domain transfer, resulting in 100K attack prompts.
To further improve the robustness and generalizability of our model against adversarial prompt attacks, we also adopt LLM-based methods for diversifying the attack prompts. This can also help prevent over-fitting on the attack template used by the works that proposed these attacks.
Moreover, in the over-refusal prompts task, annotators generate 50K questions that closely resemble potentially unsafe adversarial prompts but are deliberately crafted to be entirely safe. The primary motivation for this task is to address the over-refusal behavior commonly seen in LLMs [12], where models refuse to answer benign questions due to excessive caution. By including these prompts, we aim to train the model to better distinguish between genuinely unsafe queries and safe ones, thereby improving the model’s responsiveness while maintaining safety.
Based on 100K English (prompt, response) pairs, we translated them into Kazakh using Machine translation. Meanwhile, using the translated Kazakh prompts, we collected Kazakhstan-specific responses using GPT-4o, which results in 200K Kazakh safety instruction data and 100K English data. We combined them together to perform safety alignment.
4.2 Instruction-Tuning Setup
@@
As described in Section 4.1, each instance in our raw instruction-tuning dataset consists of a system instruction along with a user prompt and its corresponding AI-generated response. For multi-turn interactions, the data includes a sequence of multiple prompt-response pairs. Since our model builds upon Llama-3.1-Sherkala-8B, we format each raw data instance using the Llama-3.1-instruct prompt template for both supervised fine-tuning (SFT) and inference. [^6] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/ Figure 3 illustrates the transformation process from raw data to the standardized prompt format. Multiple examples are packed into sequences of context length 8192, with padding applied where necessary. As with instruction fine-tuning (IFT), the loss mask is applied only to completion/response tokens.
Instruction fine-tuning is performed over three epochs using the IFT data mix to enhance the model’s ability to generate helpful, safe, and aligned responses. The IFT dataset contains approximately 2.79 billion unique tokens. Training is conducted with a batch size of 120, a peak learning rate of 7.5e-5, and a 1% warm-up ratio, followed by cosine decay to a final learning rate of 1.5e-6.
@@ 5 Evaluation
In this section, we aim to provide a comprehensive assessment of the Sherkala (8B) model across a diverse set of evaluation dimensions, covering downstream NLP tasks, generation capabilities, and safety assessments. These evaluations are designed to rigorously measure the model’s performance and adaptability, particularly in supporting multilingual use cases across Kazakh, Russian, and English languages.
@@ 5.1 Downstream Evaluation
We conduct a comprehensive downstream evaluation, benchmarking the Sherkala (8B) model against a series of multilingual baselines that support English, Russian, and Kazakh languages. Our baseline models include general-purpose models such as BLOOM (7.1B) [65], BLOOMZ (7.1B) [48], Gemma-2 (9B) [59], Gemma-2-it (9B) [59], Qwen-2.5 (7B) [74], Qwen-2.5-Instruct (7B) [74], mGPT (13B) [67] and models from the latest Llama-3.1 series (8B) [21]. Additionally, we evaluate against LLama-3.1-KazLLM-1.0 (8B) [^7] https://huggingface.co/issai/LLama-3.1-KazLLM-1.0-8B and Irbis-v0.1, [^8] https://huggingface.co/IrbisAI/Irbis-7b-v0.1 two models fine-tuned specifically on Kazakh data.
We adopt the LM-Evaluation-Harness framework [26] to evaluate each model in a zero-shot setting, reporting accuracy for each task. Within this framework, the context string is concatenated with each candidate output string, and the answer is determined by selecting the concatenated string with the highest normalized log-likelihood.
Language Evaluation Aspect Dataset Translated? Size (set)
Kazakh World Knowledge KazMMLU [76] No 9870 (test)
MMLU [28] Yes 14042 (test)
Commonsense Reasoning Hellaswag [89] Yes 10042 (val)
PIQA [8] Yes 1838 (val)
BoolQ [16] Yes 3270 (val)
SIQA [71] Yes 1954 (val)
ARC-Challenge [9] Yes 1172 (test)
OpenBookQA [44] Yes 500 (test)
kkCOPA [46] Yes 500 (test)
Belebele [6] Yes 900 (test)
NIS MATH [46] No 100 (test)
Misinformation TruthfulQA [35] Yes 817 (val)
CrowS-Pairs [52] Yes 1508 (test)
Russian World Knowledge KazMMLU [76] No 13019 (test)
MMLU [28] Yes 10033 (test)
USE [23] No 900 (test)
WorldTree [23] Yes 525 (test)
Commonsense Reasoning OpenBookQA [44] Yes 400 (test)
English World Knowledge MMLU [28] No 14042 (test)
RACE [43] No 3498 (test)
Commonsense Reasoning Hellaswag [89] No 10042 (val)
PIQA [8] No 1838 (val)
BoolQ [16] No 3270 (val)
SIQA [71] No 1954 (val)
ARC-Challenge [9] No 1172 (test)
OpenBookQA [44] No 500 (test)
Misinformation TruthfulQA [35] No 817 (val)
CrowS-Pairs [52] No 1508 (test)

### v2

@@ 3.1 General Instruction Tuning Dataset
We curate a set of English instructions in a prompt–response pair format spanning a comprehensive range of NLP and other generative tasks. We use common datasets such as TuluV3 (Lambert et al., 2024), and also some of our data is a subset of the instruction-tuning data used for building Jais (Sengupta et al., 2023), and hence, is a combination of several publicly available datasets.
Russian:
@@
We incorporate 263K instruction-response pairs from the Grandmaster-PRO-MAX dataset, [^4] https://huggingface.co/datasets/Vikhrmodels/GrandMaster-PRO-MAX a high-quality Russian instruction-tuning corpus. This dataset covers a wide range of topics, including mathematics, coding problems, and role-playing scenarios, providing diverse linguistic and task-specific coverage for instruction fine-tuning.
3.2 Safety Alignment Dataset
@@
We developed a comprehensive English-language safety prompt collection covering eight categories of adversarial attacks and over 100 detailed safety scenarios. A team of five expert annotators initially created 1,200 high-quality “seed prompts” for direct attack alignment, drawing on prior work (Wang et al., 2024a). Building on this foundation, the expert team guided a 20-member outsourced annotation team, supported by large language models (LLMs), to generate an additional 50,000 attack prompts, ensuring diversity, linguistic relevance, and broad coverage. The dataset was further enriched with eight adversarial prompting strategies from Lin et al. (2025), which target key LLM capabilities including in-context learning, auto-regressiveness, instruction following, and domain transfer—resulting in a total of 100,000 English attack prompts.
These 100K English (prompt, response) pairs were translated into Kazakh using Google Translate. To ensure contextual appropriateness, we then used the translated Kazakh prompts to generate Kazakhstan-specific responses with GPT-4o, yielding 200K Kazakh safety instruction examples in addition to the original 100K in English. For the safety alignment stage, we subsample the dataset to 50K examples each in English and Kazakh (see Table 2).
3.3 Instruction Tuning Set-up
@@
Each instance in our raw instruction-tuning dataset consists of a system instruction along with a user prompt and its corresponding AI-generated response. For multi-turn interactions, the data includes a sequence of multiple prompt-response pairs. Since Llama-3.1-Sherkala-8B builds upon Llama-3.1, we format each raw data instance using the Llama-3.1-instruct prompt template for both supervised fine-tuning (SFT) and inference. [^5] https://www.llama.com/docs/model-cards-and-prompt-formats/meta-llama-3/ Figure 5 in Appendix G illustrates the transformation process from raw data to the standardized prompt format. Multiple examples are packed into sequences of context length 8,192, with padding applied where necessary. As with instruction fine-tuning (IFT), the loss mask is applied only to completion/response tokens. Instruction fine-tuning is performed over three epochs using the IFT data mix to enhance the model’s ability to generate helpful, safe, and aligned responses. The IFT dataset contains approximately 2.79 billion unique tokens. Training is conducted with a batch size of 120, a peak learning rate of 7.5e-5, and a 1% warm-up ratio, followed by cosine decay to a final learning rate of 1.5e-6.
@@ 4 Evaluation
We conduct three main evaluations—downstream tasks, text generation, and safety assessments—to comprehensively assess the model’s performance and adaptability. These evaluations are designed to rigorously measure its effectiveness in multilingual settings, with a particular focus on Kazakh, Russian, and English.
@@ 4.1 Downstream Evaluation
Language Evaluation Aspect Dataset Translated? Size (set)
Kazakh World Knowledge KazMMLU (Togmanov et al., 2025) No 9870 (test)
MMLU (Hendrycks et al., 2020) Yes 14042 (test)
Commonsense Reasoning Hellaswag (Zellers et al., 2019) Yes 10042 (val)
PIQA (Bisk et al., 2020) Yes 1838 (val)
BoolQ (Clark et al., 2019) Yes 3270 (val)
SIQA (Sap et al., 2019) Yes 1954 (val)
ARC-Challenge (Clark et al., 2018) Yes 1172 (test)
OpenBookQA (Mihaylov et al., 2018) Yes 500 (test)
kkCOPA (Maxutov et al., 2024) Yes 500 (test)
Belebele (Bandarkar et al., 2024) Yes 900 (test)
NIS MATH (Maxutov et al., 2024) No 100 (test)
Misinformation TruthfulQA (Lin et al., 2021a) Yes 817 (val)
CrowS-Pairs (Nangia et al., 2020) Yes 1508 (test)
Russian World Knowledge KazMMLU (Togmanov et al., 2025) No 13019 (test)
MMLU (Hendrycks et al., 2020) Yes 10033 (test)
USE (Fenogenova et al., 2024) No 900 (test)
WorldTree (Fenogenova et al., 2024) Yes 525 (test)
Commonsense Reasoning OpenBookQA (Mihaylov et al., 2018) Yes 400 (test)
English World Knowledge MMLU (Hendrycks et al., 2020) No 14042 (test)
RACE (Lai et al., 2017) No 3498 (test)
Commonsense Reasoning Hellaswag (Zellers et al., 2019) No 10042 (val)
PIQA (Bisk et al., 2020) No 1838 (val)
BoolQ (Clark et al., 2019) No 3270 (val)
SIQA (Sap et al., 2019) No 1954 (val)
ARC-Challenge (Clark et al., 2018) No 1172 (test)
OpenBookQA (Mihaylov et al., 2018) No 500 (test)
Misinformation TruthfulQA (Lin et al., 2021a) No 817 (val)
CrowS-Pairs (Nangia et al., 2020) No 1508 (test)
Datasets
To assess our model’s performance in Kazakh, Russian, and English, we use a diverse set of datasets covering world knowledge, commonsense reasoning, and misinformation detection as shown in Table 3. The evaluation in Kazakh includes MMLU (Hendrycks et al., 2020), Hellaswag (Zellers et al., 2019), BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), ARC-Challenge (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), TruthfulQA (Lin et al., 2021a), COPA (Maxutov et al., 2024), Belebele (Bandarkar et al., 2024), and CrowS-Pairs (Nangia et al., 2020), which we translated into Kazakh using Google Translate. [^6] https://translate.google.com [^7] COPA and Belebele were translated into Kazakh by the authors of the respective papers. The evaluation also incorporates two native Kazakh datasets: (1) KazMMLU (Togmanov et al., 2025), which contains 9.8K high-school-level multiple-choice questions across various subjects, and (2) NIS Math (Maxutov et al., 2024), a dataset focused on mathematical reasoning. For Russian, the evaluation consists of both translated and original datasets. Translated benchmarks include MMLU, WorldTree (Fenogenova et al., 2024), and OpenBookQA. Native datasets include KazMMLU and USE (Fenogenova et al., 2024), which are specifically designed for the Russian language. For English, we evaluate MMLU, RACE (Lai et al., 2017), Hellaswag, PIQA, BoolQ, SIQA, ARC-Challenge, OpenBookQA, TruthfulQA, and CrowS-Pairs.
@@
Our baseline models include general-purpose models such as BLOOM (7.1B) (Scao et al., 2023), BLOOMZ (7.1B) (Muennighoff et al., 2022), Gemma-2 (9B) (Riviere et al., 2024), Gemma-2-it (9B) (Riviere et al., 2024), Qwen-2.5 (7B) (Team, 2024), Qwen-2.5-Instruct (7B) (Team, 2024), mGPT (13B) (Shliazhko et al., 2024) and models from the latest Llama-3.1 series (8B) (Dubey et al., 2024). Additionally, we evaluate against LLama-3.1-KazLLM-1.0 (8B) [^8] https://huggingface.co/issai/LLama-3.1-KazLLM-1.0-8B and Irbis-v0.1, [^9] https://huggingface.co/IrbisAI/Irbis-7b-v0.1 two models fine-tuned specifically on Kazakh data.
We adopt the LM-Evaluation-Harness framework (Gao et al., 2021) to evaluate each model in a zero-shot setting, reporting accuracy for each task. Within this framework, the context string is concatenated with each candidate output string, and the answer is determined by selecting the concatenated string with the highest normalized log-likelihood.
Results for Kazakh
@@
With additional instruction fine-tuning (IFT) and safety tuning, Sherkala-Chat (8B) achieves the highest score of 47.6, making it a new benchmark for state-of-the-art Kazakh language models. This represents a +3.9+3.9 point improvement over KazLLM-1.0 (8B), a +9.9+9.9 point improvement over Irbis-v0.1 (7B), and a significant lead over all multilingual models. These results establish Sherkala-Chat (8B) as the best-performing model for Kazakh, surpassing both multilingual and Kazakh-specialized models.
Results for Russian
Table 6 in Appendix presents the evaluation results for Russian. Among multilingual models, Qwen-2.5-Instruct (7B) (38.5) and Qwen-2.5 (7B) (38.8) achieve the highest average scores, followed by Llama-3.1-KazLLM-1.0 (8B) (32.5) and Llama-3.1-Instruct (8B) (31.5). General-purpose models such as BLOOM (24.4) and Gemma-2 (21.9) perform the worst, reflecting their weaker Russian language understanding. While Qwen-2.5-Instruct remains the strongest model on Russian benchmarks, Sherkala-Chat (8B) retains the Russian capabilities of Llama-3.1 and benefits significantly from fine-tuning, narrowing the performance gap with the best multilingual models.
Results for English
As shown in Table 7 in Appendix, both Sherkala (8B) and Sherkala-Chat (8B) achieve an average score of 59.1, outperforming BLOOMZ (57.0) and LLaMA-3.1-KazLLM-1.0 (58.6), while remaining competitive with LLaMA-3.1-Instruct (60.1) and Qwen-2.5-Instruct (62.1). Although Qwen-2.5-Instruct achieves the highest average score, our model remains close to the strongest multilingual baselines. These results demonstrate that Sherkala (8B) retains strong English performance despite being primarily optimized for Kazakh.
4.2 Generation Evaluation
Datasets
We assess the text generation capabilities of Sherkala-Chat (8B) across three languages: Kazakh, Russian, and English. Following prior studies (Chiang et al., 2023; Zheng et al., 2023), we utilize the Vicuna-Instructions-80101010https://lmsys.org/blog/2023-03-30-vicuna/ and MT-Instructions-80 (Zheng et al., 2023) [^11] https://lmsys.org/blog/2023-06-22-leaderboard/ datasets and employ GPT-4o (OpenAI, 2023) as the evaluator. Both datasets were originally introduced in English; to enable evaluation in Kazakh and Russian, we translated them using Machine Translation. In addition to existing datasets, we evaluate the model on 1,000 newly constructed Kazakh-specific text generation samples, semi-automatically created using GPT-4o from Cultural Wiki and public government data (Laiyk et al., 2025). Gold responses are generated via fact extraction, followed by instruction-response pair creation and manual verification by native speakers to ensure cultural and contextual relevance. We refer to these datasets as Gov and Wiki.
Evaluation Setup
We generate model outputs using a temperature of 0.3 and a repetition penalty of 1.2 across all evaluation datasets. We compare our model against Qwen-2.5-7B-Instruct (Team, 2024), Llama-3.1-8B-Instruct (Llama, 2024), KazLLM-1.0-8B, [^12] https://huggingface.co/issai/LLama-3.1-KazLLM-1.0-8B and GPT-4o (OpenAI, 2023). Llama-3.1 serves as a strong baseline, as our model is built on it. Qwen-2.5 offers robust multilingual support, and KazLLM-1.0-8B is tailored for Kazakh, providing a competitive benchmark. GPT-4o serves as an upper bound and is also used as the evaluator, assigning scores from 0 to 10 for each model’s output. For Vicuna-Instructions-80 and MT-Instructions-80, which lack reference outputs, evaluations are based on helpfulness, relevance, accuracy, and detail. For Gov and Wiki, we additionally evaluate closeness to the reference. See the Appendix for evaluation prompts.
Results
Figure 2 shows that Sherkala-Chat (8B) consistently outperforms the baseline models in generating Kazakh text across all benchmarks. However, it still falls behind GPT-4o, which is unsurprising given the significant disparity in language model size. Notably, Sherkala-Chat (8B) surpasses the performance of KazLLM-1.0-8B, a model specifically fine-tuned for Kazakh, in the domain of Kazakh text generation. Furthermore, all models tend to achieve higher scores on MT-Instructions-80 and Vicuna-Instructions-80 compared to Gov and Wiki.
Table 8 in Appendix presents average scores and standard deviations for Kazakh, Russian, and English on the MT-Instructions-80 and Vicuna-Instructions-80 benchmarks. Sherkala-Chat (8B) shows strong performance in Kazakh and generally outperforms KazLLM-1.0-8B, but lags behind Qwen-2.5-7B-Instruct, especially in English. While it struggles with Russian in standard evaluation, pairwise comparison (Figure 4) reveals consistent wins over KazLLM-1.0-8B in both Kazakh and Russian.
4.3 Safety Evaluation
Datasets
Building on the Do-Not-Answer benchmark (Wang et al., 2024a) and its Chinese extension (Wang et al., 2024b), we construct two new safety evaluation datasets for Kazakh and Russian (Goloburda et al., 2025). These datasets are developed via translation followed by manual localization to ensure contextual relevance and linguistic quality. Additionally, we include manually crafted region-specific questions to better capture local sensitivities. Each dataset covers six risk areas, 17 harm categories, and includes approximately 4K examples. Detailed statistics are provided in Table 9, with further information on dataset construction available in Appendix F.
Rank Model Kazakh $\uparrow$ Russian $\uparrow$ English $\uparrow$
1 Claude 98.6 93.5 98.6
2 GPT-4o 98.1 87.6 95.7
3 YandexGPT 93.4 93.6 80.3
4 Llama-3.1-8B-Instruct 91.9 84.7 98.3
5 Sherkala-Chat (8B) 91.9 85.1 96.0
6 Qwen-2.5-7B-Instruct 91.4 85.1 88.1
7 Falcon3-10B-Instruct 91.3 84.7 96.8
8 KazLLM-1.0-8B 81.2 78.0 94.5
9 Aya101 78.8 84.5 96.6
10 Vikhr-Nemo-12B-Instruct 85.6 91.1
Table 5: Safety evaluation results of eight LLMs, ranked by the percentage of safe responses in the Kazakh dataset.
Evaluation Setup
Based on 939 English, 3,786 Kazakh and 4,383 Russian safety evaluation questions, we collected responses of Sherkala-Chat (8B), along with responses gathered from three commercial models (GPT-4o, Claude-3.5-Sonnet and YandexGPT), and four open-weight models: KazLLM-1.0-8B, Qwen2.5-7B-instruct, Falcon3-10B-instruct, and Llama-3.1-8B-instruct, Aya101, and Vihkr-Nemo-12B-instruct. We employ GPT-4o as a judge to assess and compare the safety rankings of eight models following prior works (Wang et al., 2024a; b). Please refer to Appendix F for further details.
Results
As shown in Table 5, models are ranked by the percentage of safe responses in the Kazakh dataset. Claude achieves the highest safety scores in both Kazakh and English, while YandexGPT is the safest for Russian. Sherkala-Chat (8B) shows improved safety over Llama-3.1-KazLLM-1.0-8B (which also builds on Llama-3.1) across all three languages. More broadly, open-weight models show relatively small differences in performance for Russian. We attribute this to adjustments in the evaluation criteria for Russian and Kazakh: rather than marking responses as unsafe based on a single violation, we allowed 1–3 violations. This change better reflects human evaluation results and helps reduce disparities across high-, medium-, and low-resource languages.
5 Related Work
Large Language Models
Transformer-based language models have revolutionized NLP, with early models such as BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019) introducing key architectural innovations and training paradigms. The emergence of larger models like GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), LLaMA-3 Touvron et al. (2023), and Falcon (Almazrouei et al., 2023) has demonstrated that scaling model size and data improves performance across a wide range of tasks. These models employ advanced techniques such as Rotary Positional Embeddings (RoPE)(Su et al., 2022) and grouped-query attention (Ainslie et al., 2023). However, they remain predominantly English-centric, limiting their effectiveness in low-resource and multilingual settings (Resnik, 2024).
Multilingual pre-training (Xue et al., 2021; Chung et al., 2023; Shliazhko et al., 2023; Scao et al., 2023; Lin et al., 2021b) supports cross-lingual transfer and wider language coverage. Models like mT5 (Xue et al., 2021), umT5 (Chung et al., 2023), mBERT (Devlin et al., 2019), XLM-RoBERTa (Conneau et al., 2020), and Aya (Üstün et al., 2024) were trained on large datasets such as mC4 (Xue et al., 2021) and Wikipedia, covering over 100 languages. However, despite including Kazakh, these models often perform worse than monolingual or specialized models, especially in tasks requiring deep linguistic understanding or domain-specific knowledge. This gap highlights the need for further research to balance broad multilingual coverage with strong performance in underrepresented languages like Kazakh.
Evaluating Large Language Models
Several benchmarks have been developed to evaluate NLP capabilities in Kazakh. KazNERD (Yeshpanov et al., 2022) is a named entity recognition dataset with over 112,000 sentences annotated across 25 entity classes. KazQAD (Yeshpanov et al., 2024) is an open-domain question-answering dataset with nearly 6,000 questions and 12,000 passage-level relevance judgments. Belebele (Bandarkar et al., 2024) includes Kazakh among its 122 languages for reading comprehension evaluation. Kardes-NLU (Senel et al., 2024) assesses cross-lingual transfer in Turkic languages, including Kazakh, across multiple NLU tasks. More recently, (Maxutov et al., 2024) systematically evaluated seven large language models on various Kazakh language tasks, highlighting the strengths and limitations of LLMs in processing Kazakh text. Despite these resources, existing Kazakh corpora do not include benchmarks for knowledge and reasoning capabilities, which are increasingly important in LLM evaluation.
6 Conclusion
We release Sherkala (8B) and Sherkala-Chat (8B) as the new state-of-the-art Kazakh-centric large language models. These open-weight models demonstrate superior performance on various Kazakh benchmarks while remaining competitive in Russian and English. Our comprehensive evaluation covers three key aspects: (1) multiple-choice question answering across knowledge, commonsense reasoning, misinformation, and bias benchmarks, (2) text generation evaluation, and (3) safety assessment. Sherkala-Chat (8B) outperforms all multilingual and Kazakh-specific models of comparable size, demonstrating its effectiveness in handling instruction-following tasks, knowledge retention, and responsible AI alignment. Beyond performance gains, we emphasize transparency by providing detailed documentation on our data collection, training process, and safety alignment strategies.
Ethics Statement
Intended Use
We believe that a number of audiences will benefit from our model:
  • Academics: those researching Kazakh natural language processing.
  • Businesses: companies targeting Kazakh-speaking audiences.
Out-of-Scope Use
Biases, Risks, and Limitations
Appendix A Training Infrastructure
The CG-2 hardware and software stack enables training extremely large models using data parallelism by relying on a special execution mode available with Cerebras Wafer Scale Clusters, called weight streaming. Weight streaming fully bypasses the complexity of 3D parallelism on traditional GPU clusters and provides simpler and higher performance scaling. You may include other additional sections here.
Appendix B Preprocessing Pipeline
Standardization
Filtering
Cleaning
Re-Filtering
Deduplication
Figure 3: Our Kazakh preprocessing pipeline.
Developing the preprocessing pipeline for Kazakh posed greater challenges compared to English. While English preprocessing pipelines benefit from numerous large-scale, open-access datasets and well-established techniques, Kazakh requires a custom-built approach. Insights gained from experiments with smaller LLMs and the preprocessing pipeline for the dataset used for Jais Sengupta et al. (2023) guided the selection of heuristics used in the final pipeline for Sherkala (8B)’s dataset. Due to the limited availability of Kazakh data, we applied less aggressive filtering than typically used for English, ensuring that valuable Kazakh content was retained.
Appendix C Prompt for Generation Evaluation
The prompt used for the evaluator is shown as follows:
Appendix D Downstream Evaluation Results for Russian and English
Results for Russian